OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±11).

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

c = 11 b = 2 c^2 = a^2 + b^2 y^2/b^2 - x^2/a^2 = 1

OpenStudy (anonymous):

is it y^2/4 -x^2/117=1

OpenStudy (anonymous):

@agent0smith